# The Dial-a-Ride Problem with Limited Pickups per Trip

### v2

@@ 3.3.6 Inequalities for fragment formulations
The above fragment-based formulations FFF and FAF may not be sufficient to lead to feasible solutions. Subtour elimination, time windows, and maximum ride time constraints will be implemented by the callbacks of infeasible path constraints, in line with Alyasiry et al. (2019); Rist & Forbes (2021). The details are as follows. As long as we encounter an incumbent solution in the branch-and-bound tree, we create chains or cycles. A chain is a sequence of fragments $(f_{1},\dots,f_{c})$ with a fragment count of $c>1$, along with $c-1$ node arcs $(a_{1},\dots,a_{c-1})$ connecting them. The load of the end node of fragment $f_{i}$ must be the same as that of the start node of $f_{i+1}$ for $11\leq i\leq c-$ 1; the start node and the end node of $a_{i}$ correspond to the end node of fragment $f_{i}$ and the start node of $f_{i+1}$, respectively, for $11\leq i\leq c-$ 1. Similarly, a cycle is a sequence of fragments $(f_{1},\dots,f_{c})$ with a fragment count of $c>1$, along with $c$ node arcs $(a_{1},\dots,a_{c})$ connecting them. A cycle has the same connection requirement as a chain but includes an additional node arc $a_{c}$ that connects $f_{c}$ back to $f_{1}$. If we obtain a chain, we connect its initial and final fragments to the origin and destination depots, respectively, to construct a path. After that, we establish a valid time schedule for the path, following the constraints of time windows and maximum ride time. If no valid schedule can be found, the chain is considered infeasible and then needs to be eliminated by constraints (35), which ensures that the total count of fragments and node arcs in such a chain does not exceed $22c-$ 2. If a cycle is detected, we then remove it using constraints (36) to ensure that the total count of its elements does not exceed $12c-$ 1, thereby achieving subtour elimination. While callbacks are a useful approach for rarely violated constraints, like the constraints mentioned above, they are not ideal for constraints that are easily violated, such as the limit on the number of pickups.
@@
In addition to the necessary constraints mentioned earlier, we incorporate three types of inequalities proposed by Rist & Forbes (2021), namely fragment-fragment, arc-fragment, and fragment-arc inequalities, and add the subset-row inequalities proposed by Jepsen et al. (2008). Fragment-fragment inequalities stipulate that when a fragment is traversed, at least one fragment from its preceding and one fragment from its succeeding fragment set must also be traversed. The preceding and succeeding fragment sets include all fragments except the ones that, when combined with the current fragment, would result in infeasible paths. Arc-fragment inequalities state that if a node arc is traversed, at least one fragment from its succeeding and one fragment from its preceding fragment set must be traversed to maintain path feasibility. Fragment-arc inequalities specify that if a fragment is traversed, at least one succeeding node arc and one preceding node arc in the corresponding sets must also be traversed. In addition, for subset-row inequalities, we follow the same approach as Rist & Forbes (2021) and Jepsen et al. (2008).
Furthermore, to enhance the formulations (both FFF and FAF), we introduce additional constraints based on the minimum number of used vehicles, as shown in constraints (37). While these constraints may be redundant, they significantly speed up the computation process. Additionally, for constraints (23), we set $w_{v}$ equal to 1 for $v\in V_{f}$.
$\sum_{f\in F}x_{vf}$ $\geq 1$ $v\in V_{f}$ (37)
@@ 3.4 MDARP-LPT: problem description and formulations
The formulations developed supra (ABF, FFF, and FAF) for DARP-LPT can also be used for MDARP-LPT, with some modifications as explained below.
For ABF, only the network flow constraints (4)-(6) need to be updated to reflect the unique origin and destination for each vehicle. The complete formulation can be found in A.
For FFF, all constraints (18)-(22) remain applicable, but the network configuration assigns unique origin and destination nodes, fragment sets, node arc sets, and node sets to each vehicle.
For FAF, in addition to the network adjustments, two further constraints, constraints (38) and (39), must be incorporated. These constraints stipulate that when a node arc starting from an origin is utilized, the subsequent fragments must also be employed, and similarly, when a node arc ending at a destination is employed, the preceding fragments must also be utilized. Specifically, for each vehicle, it possesses a unique set of node arcs that originate from its origin and a unique set of node arcs that end at its destination node.
$Y_{a}$ $\leq\sum_{f\in F^{+}_{j}}x_{fv}$ $a=(O_{v},j)\in A_{N},v\in V$ (38)
$Y_{a}$ $\leq\sum_{f\in F^{-}_{j}}x_{fv}$ $a=(j,D_{v})\in A_{N},v\in V$ (39)
Additionally, for both FFF and FAF, the model for MDARP-LPT needs to incorporate infeasible path constraints (constraints (35)) through callback procedures to handle the constraints of subtour elimination, time windows, and maximum ride time. Given that MDARP-LPT involves distinct vehicles with unique origins and destinations, constraints such as (37) based on the minimum number of vehicles are not used for MDARP-LPT.
@@ 4 Fragment generation
In this section we describe how the fragments that constitute the basis for the formulations FFF and FAF can be acquired. Section 4.1 describes the FF generation procedure of Alyasiry et al. (2019) and the RF generation method of Rist & Forbes (2021), both of which involve enumeration. The resulting fragment sets for (M)DARP-LPT are denoted as $S_{FF}$ and $S_{RF}$, respectively. From these sets, all fragments whose length exceeds $L$ are already eliminated. The set $S_{FF}$ with longer fragments tends to be larger than $S_{RF}$ with shorter fragments, yet all combinations of RFs from $S_{RF}$ can generate more (redundant) FFs than those in $S_{FF}$.
In DARP, solution paths can vary in length, potentially exceeding $L$. However, in (M)DARP-LPT, solution paths must not exceed this limit. Since $S_{FF}$ and $S_{RF}$ are primarily designed for DARP, some of their fragments (even if not exceeding $L$ in length) might only appear on solution paths that exceed $L$ and not on those that adhere to the limit (because some higher-cost paths are eliminated from consideration; see Section 4.2). Therefore, in Section 4.2, we introduce a novel path-enumerated full fragment set $S_{PFF}$ by enumerating all (M)DARP-LPT paths first and then decomposing them into full fragments. Consequently, $S_{PFF}$ must be a subset of $S_{FF}$. Another novel method in Section 4.3 is to create a mixed fragment set $S_{MF}$ by converting certain FFs from $S_{FF}$ into RFs while keeping others, aiming to leverage the strength of both FFs and RFs and potentially reduce the set size.
All these fragment sets ($S_{FF}$, $S_{RF}$, $S_{PFF}$, and $S_{MF}$) are valid for both MDARP-LPT and DARP-LPT. However, in MDARP-LPT, where vehicles have different origins and destinations, each vehicle requires a unique fragment set (though many of the fragments will still be the same for each vehicle). Therefore, the generation procedure for MDARP-LPT must be executed $|V|$ times, where $|V|$ denotes the number of vehicles. In contrast, DARP-LPT, with a fixed depot for all vehicles, only necessitates a single execution of the generation procedure.
4.1 Traditional fragment generation
Alyasiry et al. (2019) introduce FF and $S_{FF}$, while Rist & Forbes (2021) propose RF and $S_{RF}$. $S_{FF}$ exclusively contains FFs, while $S_{RF}$ exclusively contains RFs. Their intersection includes all FFs with the format “$ppppdddd$”, which can also be regarded as RFs. As previously explained, “$ppppdddd$” indicates a sequence where all pickups precede all deliveries.
@@ 4.1 Traditional fragment generation
4.2 Path-enumerated full fragment generation
Generating $S_{PFF}$ involves three steps: enumerating all paths that adhere to the problem constraints, selecting enumerated paths, and generating fragments from these selected paths. The first step is to enumerate all possible paths. Typically, such a task can be challenging due to the exponential growth in the number of paths. With limited pickups, however, this number is polynomial for $L$. To be more precise, the number of paths is bounded by $\mathcal{O}((2n)^{2L})$. The second step involves path selection. For any pair of enumerated paths that visit an identical set of locations, only the one with lower cost is kept. The third step involves extracting full fragments from the selected paths. Within each path, nodes with empty loads are identified and marked. The route paths delimited by consecutive markers are then extracted to create fragments, which are subsequently recorded in $S_{PFF}$. Notably, $S_{PFF}$ should be a subset of $S_{FF}$, as $S_{PFF}$ is derived from the selected paths, and not all FFs from $S_{FF}$ will appear on these paths. Since $S_{PFF}$ includes all FFs obtained from all non-dominated (M)DARP-LPT paths (paths with higher cost are excluded), these FFs can also be re-combined to reconstitute all non-dominated paths.
We illustrate the foregoing with an example. Consider an instance with only two customers, 1 and 2. The set $S_{FF}$ includes the following fragments: $(p1,p2,d1,d2)$, $(p1,p2,d2,\\ d1)$, $(p2,p1,d1,d2)$, $(p2,p1,d2,d1)$, $(p1,d1)$, and $(p2,d2)$. However, there are at most three useful paths: $(0,p1,d1,2n+1)$, $(0,p2,d2,2n+1)$, and then out of the four paths that include the four locations $p1$, $p2$, $d1$, and $d2$ and serve both customers, only one of the four (a shortest path) needs to be retained. As a result, $S_{PFF}$ can contain at most three fragments from the three paths and is thus clearly a subset of $S_{FF}$.
4.3 Mixed fragment generation
Path enumeration can become significantly more time-consuming as the limit on the number of pickups $L$ increases. Hence, we present an alternative procedure for creating fragments as $L$ grows. In this procedure, we initially create all FFs to get $S_{FF}$. Next, we employ a selection process to identify specific FFs to be decomposed into RFs. Finally, the remaining FFs together with the newly formed RFs constitute the set $S_{MF}$.
The first step generates FFs adhering to the (M)DARP-LPT constraints, according to the procedure described in Section 4.1, resulting in the same fragment set $S_{FF}$.
The second step decomposes all FFs into RFs and records the frequency of occurrence of each RF during this process. In a FF, pickup nodes whose preceding nodes are depots or delivery nodes, and delivery nodes whose subsequent nodes are depots or pickup nodes, are marked. The route paths between consecutive marks can be utilized to create RFs. During the process of obtaining RFs from FFs, we keep track of how frequently each RF appears in all FFs. This work is preparatory for the third step. An important aspect to note is that the resulting full RF set should be a subset of $S_{RF}$. This can be seen as follows: $S_{RF}$ encompasses all RFs obtained by partitioning FFs with the “ppppdddd” format in all possible ways. However, some RFs in $S_{RF}$ may never appear in any FF in $S_{FF}$. Specifically, these RFs might only be combined with specific other RFs to form FFs that are dominated.
The third step involves selecting FFs for decomposition, aiming to create a mixed fragment set comprising both FFs and RFs. The main objective of working with such a mixed set is to reduce the fragment count. An initial assumption is that $S_{FF}$ typically contains more fragments than $S_{RF}$. Intuitively, decomposing all FFs into RFs might seem beneficial. However, we find that if an FF includes several RFs and these RFs do not appear in any other non-dominated FFs in $S_{FF}$, decomposing it can increase the total fragment count. Furthermore, our computational results in Section 5.2.1 indicate that when RFs are used in the formulations, they may lead to weaker lower bounds compared to FFs. Based on these two observations, we only decompose a portion of FFs in $S_{FF}$. Specifically, we maintain a record of how frequently each RF appears in the FFs from the previous step. Only the FFs for which all the constituent RFs appear multiple times (at least three times, in our implementation) within all FFs in $S_{FF}$ are candidates for decomposition. Other FFs that do not meet this criterion are retained in their original form. As a result, the undecomposed FFs together with the RFs obtained through this process constitute $S_{MF}$. Thus, it can be seen that $S_{MF}$ can serve the same role as $S_{FF}$ in enumerating all (M)DARP-LPT paths and act as the input for formulations without compromising optimality.
@@ 5.1 Data generation and implementation details
To assess computational efficiency of different formulations and fragment sets, we utilize benchmark instances from Cordeau (2006), which are widely adopted in studies such as Ropke et al. (2007); Gschwind & Irnich (2015); Rist & Forbes (2021). These instances comprise two types, A and B, with type A having a tighter maximum ride time. In type A, customer demand is unit with a vehicle capacity of three, while type B involves demands from one to six customers with a vehicle capacity of six.
For DARP-LPT, we generate both basic and extended instances. The basic instances adjust the vehicle numbers $|V|$ in Cordeau’s instances. Given that DARP-LPT incorporates an additional parameter $L$ for the maximal number of pickups per trip, ensuring coverage for $n$ customers requires a minimum of $\lceil n/L\rceil$ vehicles. Here, we set $|V|$ as $\lceil n/(L-1)\rceil$, which is slightly higher than the minimum required. We also create an extra instance set to which we refer as “extended” instances, following Gschwind & Irnich, in which we increase time windows and capacity by 1/3 relative to the basic instances, thereby enhancing the diversity of the problem instances.
For MDARP-LPT, which assumes different drivers have distinct origins and destinations, we generate driver information based on Cordeau’s instances. Specifically, the first $\lceil n/\alpha\rceil$ customers are designated as drivers, whose pickup and delivery nodes becoming the origins and destinations, while the remaining customers are passengers. The value $n$ represents the number of customers in each Cordeau instance, and $\alpha$ is an auxiliary parameter. The choice $L=4$ and $\alpha=4$ implies that one driver serves three customers on average, and $\alpha=3$ has one driver per two customers on average. For drivers’ time windows, we set the earliest departure to 0, maintaining latest arrival times. Their maximum ride time is halved from the original driver’s maximum ride time (which is between six and 12 hours), as the latter is excessive for realistic scenarios such as carpooling. Unit customer demand and capacity 4 are chosen.
Table 4 summarizes our experimental design. We use $S_{PFF}$ for MDARP-LPT and $S_{MF}$ for DARP-LPT. This decision is based on the appropriateness of $S_{PFF}$ for low $L$ values, which aligns well with real carpooling scenarios in MDARP-LPT. On the other hand, $S_{MF}$ is used for comparison with $S_{RF}$ and $S_{FF}$ across diverse $L$ values, a context more suitable for DARP-LPT. We do not apply $S_{MF}$ for MDARP-LPT because each vehicle has distinct fragment sets, and mixing the FFs and RFs per vehicle would become overly complex.
@@
[b] Problem Instance $L$ $|V|$ Customer number Formulation Fragment set DARP-LPT basic $4,6$ $\lceil n/(L-1)\rceil$ $n$ ABF, FAF, FFF $S_{FF}$, $S_{RF}$, $S_{MF}$ extended MDARP-LPT basic 4 $\lceil n/\alpha\rceil$ $n-\lceil n/\alpha\rceil$ FFF $S_{FF}$, $S_{RF}$, $S_{PFF}$
@@
    Value $n$ is the number of customers in the original instance of Cordeau (2006).
@@
Since this study involves two problems (DARP-LPT and MDARP-LPT), two different fragment-based formulations, four types of fragment sets, and two types of instances (basic and extended), we need to test multiple combinations. Therefore, from the existing benchmark instances, we only choose a subset, but we ensure a variety of sizes as indicated by the instance names in all following tables.

### v3

@@ 3.3.5 Inequalities for fragment formulations
The above fragment-based formulations FFF and PSFF may not be sufficient to lead to feasible solutions. Subtour elimination, time windows, and maximum ride time constraints will be implemented by the callbacks of infeasible path constraints, in line with Alyasiry et al. [2], Rist & Forbes [18]. The details are as follows. As long as we encounter an incumbent solution in the branch-and-bound tree, we create chains or cycles. A chain is a sequence of fragments $(f_{1},\dots,f_{c})$ with a fragment count of $c>1$, along with $c-1$ node arcs $(a_{1},\dots,a_{c-1})$ connecting them. The load of the end node of fragment $f_{i}$ must be the same as that of the start node of $f_{i+1}$ for $11\leq i\leq c-$ 1; the start node and the end node of $a_{i}$ correspond to the end node of fragment $f_{i}$ and the start node of $f_{i+1}$, respectively, for $11\leq i\leq c-$ 1. Similarly, a cycle is a sequence of fragments $(f_{1},\dots,f_{c})$ with a fragment count of $c>1$, along with $c$ node arcs $(a_{1},\dots,a_{c})$ connecting them. A cycle has the same connection requirement as a chain but includes an additional node arc $a_{c}$ that connects $f_{c}$ back to $f_{1}$.
For FFF, if we obtain a chain, we connect its initial and final fragments to the origin and destination depots, respectively, to construct a path. After that, we establish a valid time schedule for the path, following the constraints of time windows and maximum ride time. If no valid schedule can be found, the chain is considered infeasible and then needs to be eliminated by constraints (32), which ensures that the total count of fragments and node arcs in such a chain does not exceed $22c-$ 2. If a cycle is detected, we then remove it using constraints (33) to ensure that the total count of its elements does not exceed $12c-$ 1, thereby achieving subtour elimination. While callbacks are a useful approach for rarely violated constraints, like the constraints mentioned above, they are not ideal for constraints that are easily violated, such as the limit on the number of pickups.
@@
For PSFF the procedure remains the same, but for the callback constraints the variables $x$ and $y$ in constraints (32) and (33) are replaced by $\sum_{l\in\mathcal{L_{P}}}X_{f_{k}l}$ and $\sum_{l\in\mathcal{L_{D}}}Y_{a_{k}l}$, respectively, and the vehicle index and vehicle set are no longer required.
3.4 Path-based formulation
Since a fragment generally spans a longer sequence than a location arc, its enumeration requires more computational effort. To further understand the trade-off, we enumerate all feasible paths and solve the problem using a path-based formulation (PBF). Given the limited number of pickups per trip, such enumeration may become practically viable, especially when the maximum number of pickups $L$ is small or when the instance size is relatively small.
The path-based formulation follows a set partitioning structure, as in Ropke et al. [20]. Let $\omega$ denote the set of all enumerated feasible paths. For each path $p\in\omega$, we define $a_{p}^{i}$ as a binary parameter that takes value 1 if location $i\in N$ is visited in path $p$, and 0 otherwise. Let $c_{p}$ represent the cost associated with path $p$. The decision variable $z_{p}$ equals 1 if path $p$ is selected, and 0 otherwise. The formulation is given below.
$\min\quad$ $\sum_{p\in\omega}c_{p}z_{p}$ (34)
$s.t.\quad$ $\sum_{p\in\omega}a_{p}^{i}z_{p}=1$ $i\in N$ (35)
$\sum_{p\in\omega}z_{p}\leq|V|$ (36)
$z_{p}\in\{0,1\}$ $p\in\omega$ (37)
The objective (34) minimizes the total cost of the selected paths. Constraints (35) ensure that each location is visited exactly once, while constraint (36) specifies the maximum number of vehicles. Finally, constraints (37) define the binary domain of the decision variables $z$.
@@ 3.5 MDARP-LPT: problem description and formulations
The formulations developed supra (ABF, FFF, PSFF, and PBF) for DARP-LPT can also be used for MDARP-LPT, with some modifications as explained below.
For ABF, only the network flow constraints (4)-(6) need to be updated to reflect the unique origin and destination for each vehicle. Specifically, the generic locations 00 and $12n+$ 1 are updated to $o_{v}$ and $d_{v}$, respectively. For the case of FFF and PSFF, all constraints remain applicable, but the network configuration assigns unique origin and destination nodes, fragment sets, node arc sets, and node sets to each vehicle. Finally, for PBF the only modification is to replace the constraint (36) by a separate inequality $\sum_{p\in\omega_{v}}z_{p}\leq 1$ for each vehicle $v$, where $\omega_{v}$ denotes the set of paths associated with vehicle $v$.
@@ 4 Fragment generation
This section describes how the fragments that constitute the basis for the formulations FFF and PSFF can be acquired. Section 4.1 outlines the enumeration-based fragment generation procedures of FFs [2] and RFs [18], resulting in fragment sets $S_{FF}$ and $S_{RF}$, respectively. Section 4.2 introduces extended fragments (EFF) and extended restricted fragments (ERF), along with their generation procedures for the corresponding sets $S_{EFF}$ and $S_{ERF}$. Finally, Section 4.3 describes a novel mixed fragment set $S_{MF}$, derived by converting FFs from $S_{FF}$ into RFs and ERFs. This approach aims to reduce the fragment count while maintaining more useful fragments.
From these sets, all fragments whose length exceeds $L$ are eliminated. The set $S_{FF}$, containing longer fragments, tends to be larger than $S_{RF}$, which contains shorter fragments. However, all combinations of RFs from $S_{RF}$ can generate more (redundant) FFs than those present in $S_{FF}$. This observation also applies to the comparison between $S_{ERF}$ and $S_{EFF}$.
All these fragment sets ($S_{FF}$, $S_{RF}$, $S_{EFF}$, $S_{ERF}$, and $S_{MF}$) are valid for both MDARP-LPT and DARP-LPT. However, in MDARP-LPT, where vehicles have different origins and destinations, each vehicle requires a unique fragment set (though many of the fragments will still be the same for each vehicle). Therefore, the generation procedure for MDARP-LPT must be executed $|V|$ times, where $|V|$ denotes the number of vehicles. In contrast, DARP-LPT, with a fixed depot for all vehicles, only necessitates a single execution of the generation procedure.
4.1 RF and FF generation
Alyasiry et al. [2] introduce FFs and $S_{FF}$, while Rist & Forbes [18] propose RFs and $S_{RF}$. $S_{FF}$ exclusively contains FFs, while $S_{RF}$ exclusively contains RFs. Their intersection is all FFs with the format “$ppppdddd$”, which can also be regarded as RFs. As previously explained, “$ppppdddd$” indicates a sequence where all pickups precede all deliveries.
@@ 4.1 RF and FF generation
4.2 EFF and ERF generation
Rist & Forbes [17] introduce the concept of EFF, which is a combination of an FF with a node arc. Similarly, we define an ERF as a combination of an RF with a node arc. To construct these EFFs and ERFs, we first generate all FFs and RFs as in Section 4.1. Then, each fragment is extended by appending either a pickup node or the destination depot. The resulting extended fragments are retained if their corresponding route paths have a feasible time schedule.
After the extension, the route path of the fragment becomes longer. The end node retains the same location information as the newly added node, while its load information remains unchanged, inheriting the load from the original fragment’s end node. Specifically, if the extended node is a pickup node rather than the destination depot, we construct an additional route path by further appending the corresponding delivery location of the extended pickup node. The extended fragment is retained only if this constructed path has a feasible schedule. Other processing steps, such as eliminating dominated fragments, remain unchanged. However, specific procedures such as combining with designated fragments for further feasibility checks are omitted, as the current enumeration process is already time-consuming and may largely exceed the actual solving time.
An additional point of attention is that when we use EFFs or ERFs, the network flow constraints and the lazy constraints in formulations FFF and PSFF need to be updated, as most node arcs are eliminated and fragments are directly connected to each other. Specifically, in FFF we replace constraints (18) and (19) with the revised constraint (38), while in PSFF we update constraints (26) and (27) to (39). The lazy constraints (32) and (33) are also adapted accordingly.
$\sum_{f\in F_{h}^{-}}x_{fv}$ $=\sum_{f\in F_{h}^{+}}x_{fv}$ $h\in P_{N},v\in V$ (38)
$\sum_{f\in F^{-}_{h}}X_{f,l+L_{f}}$ $=\sum_{f\in F^{+}_{h}}X_{fl}$ $h\in P_{N},l\in\mathcal{L_{P}}$ (39)
4.3 Mixed fragment set construction
In this subsection we introduce a new fragment set denoted as $S_{MF}$, which consists of a mixture of RFs and ERFs. This set is derived from the original full fragment set $S_{FF}$ and can be a substitute for $S_{FF}$ and $S_{ERF}$. The primary motivation for constructing such a mixed set is to reduce the overall number of fragments while preserving high-quality fragments, thereby accelerating the computation of fragment-based formulations. This new fragment set is motivated by the observation that both fragment length and count affect the quality and efficiency of fragment-based formulations. While longer fragments (e.g., FFs vs. RFs, or ERFs vs. RFs) strengthen the formulation, their corresponding set is usually larger and potentially slows down the computation, as illustrated in Section 5.
We observe that many ERFs in $S_{ERF}$ never appear in any EFF in $S_{EFF}$; in particular, some ERFs only combine with specific others to form EFFs that are ultimately dominated. Therefore, rather than exhaustively enumerating all ERFs, we propose deriving ERFs directly from $S_{EFF}$ or $S_{FF}$. Ultimately, we choose to derive ERFs from $S_{FF}$ for two main reasons. First, enumerating $S_{FF}$ requires less computational effort than $S_{EFF}$. Second, even with reduced redundancy, the number of ERFs generated from $S_{EFF}$ may still be large. Considering these points, we decide to generate fragments from $S_{FF}$. This, however, results in a mix of ERFs and RFs, rather than pure ERFs. Nevertheless, this approach significantly reduces the total number of fragments and thereby ensures computational efficiency, as demonstrated in Section 5. The generation process and the reason why mixed fragment types arise are further explained below.
To construct $S_{MF}$, we decompose each FF by identifying pickup locations whose predecessors are delivery locations. If a FF contains a number $w$ of such pickup locations, it is split into $w+1$ pieces. The first $w$ pieces are used to generate ERFs, while the final piece forms an RF. For example, consider a fragment with route path $(p_{1},p_{2},d_{2},p_{3},d_{3},p_{4},d_{4},d_{1})$. This is divided into the following three parts: $(p_{1},p_{2},d_{2},p_{3})$, $(p_{3},d_{3},p_{4})$, $(p_{4},d_{4},d_{1})$. These segments are transformed into ERFs and RFs using the method described in Section 4.1. During this transformation, it is necessary to incorporate load information at the pickup locations. For instance, the locations $p_{3}$ and $p_{4}$ are transformed into nodes $(p_{3},\{l_{1}\})$ and $(p_{4},\{l_{1}\})$, respectively. In addition, some FFs cannot be divided and are thus retained (e.g., $(p_{1},d_{1})$). Therefore, $S_{MF}$ actually includes a mix of ERFs, RFs, and FFs. Since many fragments are RFs or FFs rather than strictly ERFs or EFFs, the fragment set $S_{MF}$ maintains a relatively low overall fragment count.
In the resulting network, compared with the full fragment-based network, connections at pickup and delivery nodes with empty loads remain unchanged: fragments connect to node arcs, and node arcs connect to fragments. For pickup nodes with non-empty loads, fragments are directly connected to other fragments. Notably, there are no delivery nodes associated with non-empty loads. As a result, the network flow constraints for FFF and PSFF at nodes with empty loads remain the same as in Section 3.3.3 and Section 3.3.4, respectively, while the constraints at pickup nodes with non-empty loads are updated as described in Section 4.2.
It is worth noting that if all FFs are decomposed into RFs (instead of ERFs) and node arcs, the resulting RF set is actually a subset of $S_{RF}$. This is because $S_{RF}$ includes all RFs that can be generated from FFs (with the “ppppdddd” structure) in all possible ways. However, not all RFs in $S_{RF}$ appear in any FF from $S_{FF}$; some can only form dominated FFs when combined with others.
@@ 5.1 Data generation and implementation details
To assess computational efficiency of different formulations and fragment sets, we utilize benchmark instances from Cordeau [6], which are widely adopted in studies such as Ropke et al. [20], Gschwind & Irnich [13], Rist & Forbes [18]. These instances comprise two types, A and B, with type A having a tighter time window. In type A, customer demand is unit with a vehicle capacity of three, while type B involves demands from one to six customers with a vehicle capacity of six. In addition, we utilize benchmark instances from the SDARP problem [16, 17], which feature tighter time windows (typically three to six times the direct travel time for each customer) compared to those of Cordeau [6]. There are two main reasons for using this second dataset. First, the tighter time window constraints are more representative of our application scenarios. Second, we aim to evaluate the performance of $S_{EFF}$ for DARP-LPT, as it was shown to perform well for SDARP in Rist & Forbes [17].
For DARP-LPT, we adapt the aforementioned instances by modifying the number $|V|$ of vehicles in each instance. Given that DARP-LPT incorporates an additional parameter $L$ for the maximal number of pickups per trip, ensuring coverage for $n$ customers requires a minimum of $\lceil n/L\rceil$ vehicles. Here, we set $|V|$ as $\lceil n/(L-1)\rceil$, which is slightly higher than the minimum required.
For MDARP-LPT, which assumes different drivers have distinct origins and destinations, we generate driver information based on the aforementioned instances. Specifically, the first $\lceil n/3\rceil$ customers are designated as drivers, whose pickup and delivery nodes become the origins and destinations, while the remaining customers are passengers. The value $n$ represents the number of customers in each original instance. For drivers’ time windows, we set the earliest departure to 0, maintaining latest arrival times. Unit customer demand and capacity 4 are chosen.
Table 4 summarizes our experimental design.
@@
[b] Problem $L$ $|V|$ Cus Formulation Fragment set DARP-LPT $4,6$ $\lceil n/(L-1)\rceil$ $n$ ABF, FFF, PSFF, PBF $S_{FF}$, $S_{RF}$, $S_{EFF}$, $S_{ERF}$, $S_{MF}$ MDARP-LPT $4$ $\lceil n/3\rceil$ $\lfloor 2n/3\rfloor$ PSFF, PBF $S_{FF}$, $S_{RF}$, $S_{MF}$
@@
    Cus = customer number; Value $n$ is the number of customers in the original instance.
@@
Since this study involves two problems (DARP-LPT and MDARP-LPT), four different formulations, and five types of fragment sets, we need to test multiple combinations. Therefore, from the benchmark instances provided by Cordeau [6], we only choose a subset, but we ensure a variety of sizes as indicated by the instance names in all following tables. As the instances from Riedler & Raidl [16] have tighter time windows, which adheres more closely to the carpooling and crowdshipping scenarios for the problem addressed in this study, we retain all their instances.